# A Gaussian process framework for testing general relativity with gravitational waves

### v1

@@ Abstract
Black holes (162) — Compact objects (288) — Gravitational wave astronomy (675) — Gravitational waves (678)
1 Introduction
@@
Inspiral-merger-ringdown consistency tests check if the early and late stages of binary evolution are consistent (see, e.g., Hughes & Menou, 2005; Ghosh et al., 2017; Abbott et al., 2016a; Abbott et al., 2019, 2021a, 2021b). Parameterised post-Newtonian tests are performed by introducing deviations in the post-Newtonian waveform coefficients predicted for GR (see, e.g., Abbott et al., 2016a; Abbott et al., 2019, 2021a, 2021b). In residual strain tests, several flexible models have been used to search for unmodelled, coherent excess power in gravitational-wave residual data (after the best-fit GR template has been subtracted). These include spline models, which consist of smoothly-connected piecewise polynomials (Edelman et al., 2021), and algorithms such as BayesWave (Cornish & Littenberg, 2015, see e.g., Abbott et al. (2016a); Abbott et al. (2019, 2021a, 2021b)), which models signals as a sum of flexible wavelets and allows the number of wavelets to vary during sampling.
@@
We structure this paper as follows. In Section 2, we describe our Gaussian process framework. In Section 3 we demonstrate this framework with simulated data. In Section 4, we apply our method to the 60 binary black hole gravitational-wave events from the third Gravitational-Wave Transient Catalogue (GWTC-3, Abbott et al., 2023) detected by both the LIGO Hanford (H1) and Livingston (L1) observatories. We find no evidence for deviations from GR, and limit the fractional deviation in gravitational-wave strain to as low as $7\%$ (90% credibility) of the strain of GW190701_203306.
2 Formalism
2.1 The Gaussian process likelihood
Our strain data $h$ consists of three components: a signal predicted by GR $s_{\text{GR}}$, instrumental noise $n$, and a signal from physics beyond GR $\delta s$: [^1] In practice, $\delta s$ also contains unmodelled effects such as noise and waveform systematics—see Gupta et al. (2024) for a recent review. We discuss this further in Section 5.
@@ 2.1 The Gaussian process likelihood
where $P(f_{j})$ is the single-sided noise power spectral density (Thrane & Talbot, 2019a), [^2] In practice, the data are windowed with some function $w(t)$, which reduces the noise power spectral density by a factor of $\overline{w^{2}}$. $\Delta f$ is the frequency-bin width and $\delta_{ij}$ is a Kronecker delta. Here and throughout, repeated indices do not imply summation.
@@
Given this assumption, deviations from GR are described probabilistically rather than with parameterised waveforms (e.g., Abbott et al., 2016b, 2021c). The form of S is determined by our prior beliefs about deviations from GR. In general, S may depend on parameters that we denote $\Lambda$.
2.2 Kernel design
To derive an expression for S, we first design a kernel K, which describes the covariance between $\delta s(t)$ at different times. [^3] Here, we implicitly assume that the deviation from GR produces a GW that is +,×+$\times$+ , × polarised. The more general case, where the deviation appears as a non-standard polarisation mode, is discussed in the companion paper Cheung et al. (2025). In choosing our kernel, we adopt the following prior beliefs about $\delta s(t)$:
@@ 2.2 Kernel design
  • The deviation is localised in time so that $\delta s(t)$ goes to zero before and after the merger on a time scale comparable to the inverse characteristic frequency $1/f_{0}$.
@@
is the absolute value of the difference of two sample times $t_{i}\text{ and }t_{j}$. In Fig. 1, we plot draws of $\delta s(t)$ from the kernel (Eq. 2.2). These random draws showcase that this kernel produces the features consistent with our prior beliefs. The kernel is composed of several commonly used kernel functions, which together enforce our prior beliefs:
@@
Figure 1: Example deviations from GR drawn from our kernel (Eq. 2.2), with scale factor $k_{0}=10^{-45}$, characteristic frequency $f_{0}=100\,\mathrm{Hz}$, width $w=2.0$, and length $l=2.5$.
Since the noise in gravitational-wave observatories is described in the frequency domain, the kernel from Eq. 2.2 must be transformed before it can be incorporated with the detector noise in the likelihood function. The frequency-domain kernel is
@@
2.3 Detector response
Given our assumption that the deviation from GR consists of $+$ and $\times$ polarised GWs, the strain measured in detector $\mu$ is
@@ 2.3 Detector response
[…]
@@
where $\tau_{\mu\nu}$ is the time delay between detectors $\mu$ and $\nu$, which implicitly depends on the sky location of the source. We assume that the $+$ and $\times$ modes are uncorrelated, [^4] One is free to adopt other assumptions, for example, that the source is linearly polarised. such that
@@
We assume that the $+$ and $\times$ modes behave the same way on average, such that
@@
We assume each realisation of $\delta s_{+}(f_{i})$ and $\delta s_{\times}(f_{j})$ are independent draws from a normal distribution with covariance $\textbf{{K}}_{ij}$, which depends on parameters $\Lambda\equiv\{k_{0},w,f_{0},l\}$. Given our assumptions, the signal covariance matrix S (defined in Eq. 4) can be written in terms of the kernel K and the antenna factors for two detectors as a block matrix: [^5] We derive the form of S for $>2$ detectors in Appendix A. {widetext}
@@ 2.4 Building the likelihood
Since $s_{\text{GR}}$ depends on the binary parameters $\theta$, $\delta h$ also depends on $\theta$. For simplicity, we assume below that the maximum-likelihood estimate of $s_{\text{GR}}$ perfectly subtracts the GR signal from the data. However, one can in principle marginalise over uncertainty in $\theta$ with importance sampling. We demonstrate this marginalisation in Section 5. We find that this has a minimal effect on our confidence in detections of deviations from GR. Therefore, we will proceed using the max-likelihood estimate of $s_{\text{GR}}$.
@@
${\cal L}(\delta h|\Lambda)=$ $\frac{1}{2\pi\det\textbf{{C}}(\Lambda)}\exp\left(-\frac{1}{2}|\delta h^{\dagger}\textbf{{C}}^{-1}(\Lambda)\delta h|\right),$ (14)
where $\textbf{{C}}(\Lambda)$ is the total covariance matrix
@@
3 Demonstration
3.1 Tests on simulated signals
We first analyse a simulated signal drawn from the signal covariance matrix S (Eq. 2.3) with hyper-parameter values $k_{0}=2.0\times 10^{-43}$, $f_{0}=100\,\mathrm{Hz}$, $w=2.0$, and $l=2.5$. We add this signal to a randomly selected $2\,\mathrm{s}$ long data segment (beginning at GPS = $1267702309\,\mathrm{s}$) from the LIGO-Virgo-KAGRA collaboration’s third observing run (O3), during which both the LIGO H1 and L1 observatories were in observing mode and no GW signal is evident. The injection has an optimal network SNR of $9.4$. We use gravitational-wave strain data from the Gravitational Wave Open Science Centre (GWOSC; Abbott et al., 2023), and perform Gaussian process regression using the nested sampler dynesty (Speagle, 2020) as implemented in Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020). We construct the noise power spectral density (PSD, Eq. 2) using the $64\,\mathrm{s}$ of data prior to the GPS trigger time. We split the data into overlapping segments and use the median to estimate the PSD.
We choose priors on the Gaussian process hyper-parameters as follows: $f_{0}$ is distributed uniformly on the interval $[50\,\mathrm{Hz},560\,\mathrm{Hz}]$, as we expect this range to capture a deviation from GR arising in a GW150914-like event, as discussed in Subsection 2.2. On average, signals with characteristic frequencies above $560\,\mathrm{Hz}$, with other parameters drawn from their respective priors, have more than $90\%$ of their SNR above the maximum observing frequency of $512\,\mathrm{Hz}$. Our priors on $w$ and $l$ are log-uniformly distributed on the interval $[0.1,5]$. The lower bound of these priors is chosen to include signals in which approximately half a cycle is visible in the time domain strain of typical signals drawn from the kernel. The upper bound on these parameters is chosen so that a maximum of approximately 10 to 20 cycles are visible. Finally, $k_{0}$ is log-uniform distributed over the interval $[10^{-46},10^{-41}]$. The minimum bound is chosen to correspond to a signal that is indistinguishable from zero; the average optimal network signal-to-noise ratio for a signal with $k_{0}=10^{-46}$, with other parameters drawn from their associated priors, is approximately $0.10$. The maximum bound is chosen to correspond to an unambiguous detection; the average optimal network signal-to-noise ratio for a signal with $k_{0}=10^{-41}$, with other parameters drawn from their associated priors, is approximately $30$. [^6] These numbers are obtained by analysing a 2 s long segment of typical noise beginning at GPS trigger time $1267702309\,\mathrm{s}$. We choose a log-uniform distributed prior for $w$, $l$ and $k_{0}$ as we do not know the typical orders of magnitude of the length scale, or amplitude, a deviation from GR will manifest as.
@@ 3.1 Tests on simulated signals
[…]
which is the ratio of the Bayesian evidence for $k_{0}>0$ to the Bayesian evidence for $k_{0}=0$, sometimes called the noise evidence. If $\ln\mathcal{B}\ll 0$, the noise model is preferred over the signal model, and if $\ln{\cal B}\gg 0$, the signal model is preferred. Using Eq. 16, we obtain $\ln{\mathcal{B}}=21.9$ in support of the signal hypothesis for the optimal network SNR $=9.3$ signal injection, showing that we correctly find evidence for a deviation.
Figure 2: Posterior distribution of the Gaussian process hyper-parameters for a simulated signal with a deviation from GR drawn from our kernel. The contours are the $1$,$2$ and $3\sigma$ intervals. The true values are plotted as orange cross hairs, and are included within at least the $3\sigma$ interval for all hyper-parameters.
@@ 3.2 Tests on noise
We check that these data segments do not contain previously identified glitches by checking for the presence of either CAT1 or CAT2 data-quality vetoes (see, e.g. Abbott et al., 2023). We also check for coincidence of any data segments with glitches identified by the GravitySpy framework (Zevin et al., 2017; Glanzer et al., 2023) with higher than 90% confidence. We do not include a data segment if its center lies within 1.5 s of such a glitch.
To ensure the analysis of noise segments is as similar as possible to the analysis of real GW events, we inject a GW150914-like gravitational-wave signal into each segment. We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform. We use standard priors on the right ascension $\alpha$, declination $\delta$, inclination angle $\theta_{JN}$, azimuthal angle $\phi_{12}$, coalescence phase $\phi_{c}$, polarisation angle $\psi$, geocent time $t_{0}$, and spin tilts $\theta_{1}$ and $\theta_{2}$. We choose a uniform prior on the spin magnitudes $a_{1}$ and $a_{2}$ in the range $[0,0.99]$. We choose a uniform prior in comoving luminosity distance in the range $[10,10^{4}]\,\mathrm{Mpc}$. We constrain $m_{1}$ and $m_{2}$ in the range $[10,200]\,\mathrm{M_{\odot}}$, use a uniform prior in chirp mass $\mathcal{M}{}$ in the range $[20,60]\,\mathrm{M_{\odot}}$ and a uniform prior in mass ratio $q$ in the range $[0.05,1]$. We marginalise over time, distance and phase (see, e.g. Thrane & Talbot, 2019b) during sampling. We form residual data by subtracting the best-fit (maximum-likelihood) GR template from the data. We perform Gaussian process regression on the remaining 174 signals using the procedure detailed in the preceding section.
@@
3.3 Tests with other signals
To test the flexibility of our formalism, we analyse a GR-violating signal not drawn from our kernel. In particular, we consider a GR-violating waveform where the signal is described by IMRPhenomPv2 (Schmidt et al., 2012), except the intermediate-stage post-Newtonian phase coefficient $\beta_{2}$ is changed from the value predicted by GR. This parameter is particularly important for short-duration, binary black hole mergers (Abbott et al., 2016a). Using the Bilby TGR package (Ashton et al., 2024), we inject a gravitational-wave signal with GW150914-like parameters, but we change $\beta_{2}$ by $\delta\beta_{2}=1.0$.
We inject the gravitational-wave signal into $1\,\mathrm{s}$ of simulated Gaussian noise, coloured according to the LIGO design sensitivity. We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform. We use the same binary-parameter priors as in Section 3.2. We form residual data by subtracting the best-fit (maximum-likelihood) GR template from the data. This residual data necessarily contains a deviation from GR (due to the imperfect subtraction of a GR signal from a non-GR signal), with an optimal network SNR $=12.1$. We center the data at the max-likelihood merger time of the event. We perform Gaussian process regression on this segment, with the same procedure as before.
In Fig. 5, we show the posterior distribution for our kernel hyper-parameters. We see that the minimum value of $k_{0}=10^{-46}$ is excluded from the $90\%$ credible interval. We plot the whitened deviation-strain reconstruction in Fig. 6, from which it is clear a deviation from GR is present. The signal-to-noise Bayes factor for this event is $\ln{\mathcal{B}}=13.02$, with an associated $p$-value of $p=0.005$ when compared to the background distribution, suggesting strong statistical significance. This demonstrates our model is flexible enough to capture deviations from GR that are not drawn from the kernel.
Figure 5: Posterior distribution of the Gaussian process hyper-parameters for a simulated gravitational-wave signal with a parameterised deviation from GR in the form of a $\delta\beta_{2}\neq 0$ phase coefficient. The contours are the $1$,$2$ and $3\sigma$ intervals.

### v2

@@ Abstract
I Introduction
@@
Inspiral-merger-ringdown consistency tests check if the early and late stages of binary evolution are consistent (see, e.g., Hughes and Menou, 2005; Ghosh et al., 2017; Abbott and others, 2016b, 2019, 2021b, 2021c). Parameterised post-Newtonian tests are performed by introducing deviations in the post-Newtonian waveform coefficients predicted for GR (see, e.g., Abbott and others, 2016b, 2019, 2021b, 2021c). In residual strain tests, several flexible models have been used to search for unmodelled, coherent excess power in gravitational-wave residual data (after the best-fit GR template has been subtracted). These include spline models, which consist of smoothly-connected piecewise polynomials (Edelman et al., 2021), and algorithms such as BayesWave (Cornish and Littenberg, 2015, see e.g., Abbott and others (2016b, 2019, 2021b, 2021c)), which models signals as a sum of flexible wavelets and allows the number of wavelets to vary during sampling.
@@
We structure this paper as follows. In Section II, we describe our Gaussian process framework. In Section III we demonstrate this framework with simulated data. In Section IV, we apply our method to the 60 binary black hole gravitational-wave events from the third Gravitational-Wave Transient Catalogue (GWTC-3, Abbott et al., 2023) detected by both the LIGO Hanford (H1) and Livingston (L1) observatories. We find no evidence for deviations from GR, and limit the fractional deviation in gravitational-wave strain to as low as $7\%$ (90% credibility) of the strain of GW190701_203306.
II Formalism
II.1 The Gaussian process likelihood
Our strain data $h$ consists of three components: a signal predicted by GR $s_{\text{GR}}$, instrumental noise $n$, and a signal from physics beyond GR $\delta s$: [^1] In practice, $\delta s$ also contains unmodelled effects such as noise and waveform systematics—see Gupta et al. (2024) for a recent review. We discuss this further in Section V.
@@ II.1 The Gaussian process likelihood
where $P(f_{j})$ is the single-sided noise power spectral density (Thrane and Talbot, 2019a), [^2] In practice, the data are windowed with some function $w(t)$, which reduces the noise power spectral density by a factor of $\overline{w^{2}}$. $\Delta f$ is the frequency-bin width and $\delta_{ij}$ is a Kronecker delta. Here and throughout, repeated indices do not imply summation.
@@
Given this assumption, deviations from GR are described probabilistically rather than with parameterised waveforms (e.g., Abbott and others, 2016a, 2021a). The form of S is determined by our prior beliefs about deviations from GR. In general, S may depend on parameters that we denote $\Lambda$.
II.2 Kernel design
To derive an expression for S, we first design a kernel K, which describes the covariance between $\delta s(t)$ at different times. [^3] Here, we implicitly assume that the deviation from GR produces a GW that is $\times$ polarised. The more general case, where the deviation appears as a non-standard polarisation mode, is discussed in the companion paper Cheung et al. (2025). In choosing our kernel, we adopt the following prior beliefs about $\delta s(t)$:
@@ II.2 Kernel design
  • The deviation is localised in time so that $\delta s(t)$ goes to zero before and after the merger on a time scale comparable to the inverse characteristic frequency $1/f_{0}$
@@
is the absolute value of the difference of two sample times $t_{i}\text{ and }t_{j}$. In Fig. 1, we plot draws of $\delta s(t)$ from the kernel (Eq. II.2). These random draws showcase that this kernel produces the features consistent with our prior beliefs. The kernel is composed of several commonly used kernel functions, which together enforce our prior beliefs:
@@
Figure 1: Example deviations from GR drawn from our kernel (Eq. II.2), with scale factor $k_{0}=10^{-45}$, characteristic frequency $f_{0}=100\,\mathrm{Hz}$, width $w=2.0$, and length $l=2.5$.
Since the noise in gravitational-wave observatories is described in the frequency domain, the kernel from Eq. II.2 must be transformed before it can be incorporated with the detector noise in the likelihood function. The frequency-domain kernel is
@@
II.3 Detector response
Given our assumption that the deviation from GR consists of ++ and $\times$ polarised GWs, the strain measured in detector $\mu$ is
@@ II.3 Detector response
Here, $F_{+,\times}$ are the antenna response functions, which depend on the sky location $(\text{ra},\text{dec})$ and polarization angle $({\psi})$ (Nishizawa et al., 2009). The signal in detector $\nu$ is related to the signal in detector $\mu$ by a phase shift
@@
where $\tau_{\mu\nu}$ is the time delay between detectors $\mu$ and $\nu$, which implicitly depends on the sky location of the source. We assume that the ++ and $\times$ modes are uncorrelated, [^4] One is free to adopt other assumptions, for example, that the source is linearly polarised. such that
@@
We assume that the ++ and $\times$ modes behave the same way on average, such that
@@
We assume each realisation of $\delta s_{+}(f_{i})$ and $\delta s_{\times}(f_{j})$ are independent draws from a normal distribution with covariance $\textbf{{K}}_{ij}$, which depends on parameters $\Lambda\equiv\{k_{0},w,f_{0},l\}$. Given our assumptions, the signal covariance matrix S (defined in Eq. 4) can be written in terms of the kernel K and the antenna factors for two detectors as a block matrix: [^5] We derive the form of S for >2>2 detectors in Appendix A.
@@ II.4 Building the likelihood
Since $s_{\text{GR}}$ depends on the binary parameters $\theta$, $\delta h$ also depends on $\theta$. For simplicity, we assume below that the maximum-likelihood estimate of $s_{\text{GR}}$ perfectly subtracts the GR signal from the data. However, one can in principle marginalise over uncertainty in $\theta$ with importance sampling. We demonstrate this marginalisation in Section V. We find that this has a minimal effect on our confidence in detections of deviations from GR. Therefore, we will proceed using the max-likelihood estimate of $s_{\text{GR}}$.
@@
${\cal L}(\delta h|\Lambda)=$ $\frac{1}{\pi^{n}\det\textbf{{C}}(\Lambda)}\exp\left(-\delta h^{\dagger}\textbf{{C}}^{-1}(\Lambda)\delta h\right),$ (14)
where $n$ is the number of frequency bins and $\textbf{{C}}(\Lambda)$ is the total covariance matrix
@@
III Demonstration
III.1 Tests on simulated signals
We first analyse a simulated signal drawn from the signal covariance matrix S (Eq. II.3) with hyper-parameter values $k_{0}=2.0\times 10^{-43}$, $f_{0}=100\,\mathrm{Hz}$, $w=2.0$, and $l=2.5$. We add this signal to a randomly selected $2\,\mathrm{s}$ long data segment (beginning at GPS = $1267702309\,\mathrm{s}$) from the LIGO-Virgo-KAGRA collaboration’s third observing run (O3), during which both the LIGO H1 and L1 observatories were in observing mode and no GW signal is evident. The injection has an optimal network signal-to-noise ratio (SNR) of $9.4$. We use gravitational-wave strain data from the Gravitational Wave Open Science Centre (GWOSC; Abbott and others, 2023), and perform Gaussian process regression using the nested sampler dynesty (Speagle, 2020) as implemented in Bilby (Ashton et al., 2019; Romero-Shaw et al., 2020). We construct the noise power spectral density (PSD, Eq. 2) using the $64\,\mathrm{s}$ of data prior to the GPS trigger time. We split the data into overlapping segments and use the median to estimate the PSD.
We choose priors on the Gaussian process hyper-parameters as follows: $f_{0}$ is distributed uniformly on the interval $50\,\mathrm{Hz},560\,\mathrm{Hz}]$, as we expect this range to capture a deviation from GR arising in a GW150914-like event, as discussed in Subsection II.2. On average, signals with characteristic frequencies above $560\,\mathrm{Hz}$, with other parameters drawn from their respective priors, have more than $90\%$ of their SNR above the maximum observing frequency of $512\,\mathrm{Hz}$. Our priors on $w$ and $l$ are log-uniformly distributed on the interval $[0.1,5]$. The lower bound of these priors is chosen to include signals in which approximately half a cycle is visible in the time domain strain of typical signals drawn from the kernel. The upper bound on these parameters is chosen so that a maximum of approximately 10 to 20 cycles are visible. Finally, $k_{0}$ is log-uniform distributed over the interval $10^{-46},10^{-41}]$. The minimum bound is chosen to correspond to a signal that is indistinguishable from zero; the average optimal network signal-to-noise ratio for a signal with $k_{0}=10^{-46}$, with other parameters drawn from their associated priors, is approximately $0.10$. The maximum bound is chosen to correspond to an unambiguous detection; the average optimal network signal-to-noise ratio for a signal with $k_{0}=10^{-41}$, with other parameters drawn from their associated priors, is approximately 3030.666These numbers are obtained by analysing a 2 s long segment of typical noise beginning at GPS trigger time $1267702309\,\mathrm{s}$. We choose a log-uniform distributed prior for $w$, $l$ and $k_{0}$ as we do not know the typical orders of magnitude of the length scale, or amplitude, a deviation from GR will manifest as.
@@ III.1 Tests on simulated signals
$\mathcal{B}=\frac{\int dk_{0}\,\mathcal{L}(\delta h|k_{0})\,\pi(k_{0})}{\mathcal{L}(\delta h|k_{0}=0)\ },$ (16)
which is the ratio of the Bayesian evidence for $k_{0}>0$ to the Bayesian evidence for $k_{0}=0$, sometimes called the noise evidence. If $\ln\mathcal{B}\ll 0$, the noise model is preferred over the signal model, and if $\ln{\cal B}\gg 0$, the signal model is preferred. Using Eq. 16, we obtain $\ln{\mathcal{B}}=21.9$ in support of the signal hypothesis for the optimal network SNR =9.3=9.3 signal injection, showing that we correctly find evidence for a deviation.
Figure 2: Posterior distribution of the Gaussian process hyper-parameters for a simulated signal with a deviation from GR drawn from our kernel. The contours are the 11,22 and $3\sigma$ intervals. The true values are plotted as orange cross hairs, and are included within at least the $3\sigma$ interval for all hyper-parameters.
@@ III.2 Tests on noise
We check that these data segments do not contain previously identified glitches by checking for the presence of either CAT1 or CAT2 data-quality vetoes (see, e.g. Abbott and others, 2023). We also check for coincidence of any data segments with glitches identified by the GravitySpy framework (Zevin et al., 2017; Glanzer and others, 2023) with higher than 90% confidence. We do not include a data segment if its center lies within 1.5 s of such a glitch.
To ensure the analysis of noise segments is as similar as possible to the analysis of real GW events, we inject a GW150914-like gravitational-wave signal into each segment. We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform. We use standard priors on the right ascension $\alpha$, declination $\delta$, inclination angle $\theta_{JN}$, azimuthal angle $\phi_{12}$, coalescence phase $\phi_{c}$, polarisation angle $\psi$, geocent time $t_{0}$, and spin tilts $\theta_{1}$ and $\theta_{2}$. We choose a uniform prior on the spin magnitudes $a_{1}$ and $a_{2}$ in the range $[0,0.99]$. We choose a uniform prior in comoving luminosity distance in the range $10,10^{4}]\,\mathrm{Mpc}$. We constrain $m_{1}$ and $m_{2}$ in the range $10,200]\,\mathrm{M_{\odot}}$, use a uniform prior in chirp mass $\mathcal{M}{}$ in the range $20,60]\,\mathrm{M_{\odot}}$ and a uniform prior in mass ratio $q$ in the range $[0.05,1]$. We marginalise over time, distance and phase (see, e.g. Thrane and Talbot, 2019b) during sampling. We form residual data by subtracting the best-fit (maximum-likelihood) GR template from the data. We perform Gaussian process regression on the remaining 174 signals using the procedure detailed in the preceding section.
@@
III.3 Tests with other signals
To test the flexibility of our formalism, we analyse a GR-violating signal not drawn from our kernel. In particular, we consider a GR-violating waveform where the signal is described by IMRPhenomPv2 (Schmidt et al., 2012), except the intermediate-stage post-Newtonian phase coefficient $\beta_{2}$ is changed from the value predicted by GR. This parameter is particularly important for short-duration, binary black hole mergers (Abbott and others, 2016b). Using the Bilby TGR package (Ashton et al., 2024), we inject a gravitational-wave signal with GW150914-like parameters, but we change $\beta_{2}$ by $\delta\beta_{2}=1.0$.
We inject the gravitational-wave signal into $1\,\mathrm{s}$ of simulated Gaussian noise, coloured according to the LIGO design sensitivity. We perform parameter estimation to obtain posteriors on the binary parameters of the IMRPhenomPv2 waveform. We use the same binary-parameter priors as in Section III.2. We form residual data by subtracting the best-fit (maximum-likelihood) GR template from the data. This residual data necessarily contains a deviation from GR (due to the imperfect subtraction of a GR signal from a non-GR signal), with an optimal network SNR =12.1=12.1. We center the data at the max-likelihood merger time of the event. We perform Gaussian process regression on this segment, with the same procedure as before.
In Fig. 5, we show the posterior distribution for our kernel hyper-parameters. We see that the minimum value of $k_{0}=10^{-46}$ is excluded from the $90\%$ credible interval. We plot the whitened deviation-strain reconstruction in Fig. 6, from which it is clear a deviation from GR is present. The signal-to-noise Bayes factor for this event is $\ln{\mathcal{B}}=13.02$, with an associated $p$-value of $p=0.005$ when compared to the background distribution, suggesting strong statistical significance. The signal is better fit for times $t>0$, as this part of the signal is better represented using a single characteristic frequency, which matches the model, compared to the time-dependent frequency evolution for $t<0$. One could better model such ‘chirping’ deviations using a kernel with multiple characteristic frequencies, or a time-dependent characteristic frequency. It also has higher frequency content than the signal at $t<0$, and since the noise at lower frequencies is greater, this part of the signal is higher SNR, and therefore better captured by the model. This demonstrates our model is flexible enough to capture deviations from GR that are not drawn from the kernel.
Figure 5: Posterior distribution of the Gaussian process hyper-parameters for a simulated gravitational-wave signal with a parameterised deviation from GR in the form of a $\delta\beta_{2}\neq 0$ phase coefficient. The contours are the 11,22 and $3\sigma$ intervals.